OpenStudy (anonymous):

11. What is the interest earned when $1,500 is invested at a rate of 2.5% for 3 years? A. $1,612.50 B. $112.50 C. $11,250 D. $1,387.50 12. What is the interest earned when $7,000 is invested for 6 months at a rate of 4.25% ? A. $148.75 B. $7,148.75 C. $1,785.00 D. $178,500.00 13. What is the BALANCE of an account after 24 months when $3,700 is invested at 5% ? A. $370.00 B. $4,440.00 C. $4,070.00 D. $444,000.00

OpenStudy (anonymous):

11. multiply 1500 by 2.5% to see how much your cash went up for 1 year 1500(1.025) = 1537.5 to find interest subrtract that number by 1500 so see how much it increase 1537.5-1500 =37.5 then multiply by 3 to get it for 3 yrs 37.5(3) = 112.5 <- thats your interest 12. same thing but 6 months 7000(1.0425) = 7297.5 7297.5 - 7000 = 297.5 <- interest for 1 year 297.5(6/12) = 148.75 <- interest for 6 months 13. same thing but its 2 years 3700(1.05) = 3885 3885 - 3700 = 185 <- interest for 1 year 185(2) = 370 <- interst for 2 years 3700 + 370 = 4070 <- balance after 2 years

OpenStudy (anonymous):

thank you